OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)?

OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

can some one help

OpenStudy (camerondoherty):

Slope = \[\Large\frac{ Rise }{ Run }\] So i Graphed the points: http://prntscr.com/4rvho0 If you look at the rise it is 2 and the run is one 2/1 = 2 So your slope would be -2, Since it is decreasing Another way to determine the slope is by using the point slop formula \[m=\frac{ (y_1-y_2) }{ (x_1-x_2) }\] So if \(y_1\) = 3 \(y_2\) = -7 \(x_1\) = -1 \(x_2\) = 4 Then you substitute and solve \[m=\frac{ (3+7) }{ (-1-4) }\] Which is \(\Large\frac{ 10 }{ -5 }\) Can you solve That? It ends up being -2 So your slope would be -2 c:

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wrong

OpenStudy (camerondoherty):

How is it wrong?

OpenStudy (camerondoherty):

Please Explain

OpenStudy (anonymous):

i 5;'l87deberruyfelioy77808ufkt3

OpenStudy (anonymous):

??????????????

OpenStudy (camerondoherty):

Joy, Its Not wrong Hope I helped c:

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

wwwwwwwwwwwwrrrrrrrrrrrrrrrrrrroooooooooooooonnnnnnnnnngggggggggggggg

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

don't be childish @edmodo

OpenStudy (anonymous):

if u know what is it

OpenStudy (anonymous):

@camerondoherty can u help with more

OpenStudy (anonymous):

no